OpenStudy (anonymous):

I need help on my algebra Factor: x2 - 100

OpenStudy (anonymous):

hint a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

@deer1214 try to solve it and i'll check ur answer

OpenStudy (anonymous):

ok so its x^2-100^2

OpenStudy (anonymous):

yeah factor it for example x^2-9=(x-3)(x+3)

OpenStudy (anonymous):

so it would be x^2-10=(x-10)(x+10)

OpenStudy (anonymous):

yup exactly :)

OpenStudy (anonymous):

but u made a typo it's x^2-100

OpenStudy (anonymous):

so thats the answer to my question

OpenStudy (anonymous):

x^2-100=(x-10)(x+10)

OpenStudy (anonymous):

ok so its x^2-100=(x-10)(x+10)

OpenStudy (anonymous):

yup :D

OpenStudy (anonymous):

thank you som much

OpenStudy (anonymous):

u r welcome :D

OpenStudy (anonymous):

medal ?? if u don't mind :D

OpenStudy (anonymous):

thanx :)

OpenStudy (anonymous):

subtract 100 and just square the x and get (x+10)(x-10)

OpenStudy (anonymous):

@31356

OpenStudy (31356):

Correct! :D